OpenStudy (anonymous):

the sum of two consecutive odd integers is -188. find the integers

OpenStudy (anonymous):

do you know what "consecutive" means in this case?

OpenStudy (anonymous):

Integers that follow in sequence

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yeah and they are both odd so i would guess and check for example \(-95-97=-192\) is not right try another pair close

OpenStudy (anonymous):

you could also solve the equation \(n+n+2=-188\) or \(2n+2=-188\) if you like to solve equations instead of guessing if it was me i would guess and check

OpenStudy (anonymous):

can they both be -95?

OpenStudy (anonymous):

actually. -95 and =93

OpenStudy (anonymous):

no they have to be different when you solve \[2n+2=-188\\2n=-190\\n=-95\] then you just solved for \(n\) the other one is \(n+2=-95+2=-93\)

OpenStudy (anonymous):

yes that is it